OpenStudy (anonymous):

Simplify - 3/5 dived by 7/6

OpenStudy (anonymous):

multiple choice - 7/ 10 ,- 18 / 35 , 18 / 35 , 7/10

OpenStudy (tal_tal):

Thank you for putting the answers :)

OpenStudy (anonymous):

lol no prob

OpenStudy (texaschic101):

when dividing fractions, flip what you are dividing by, then multiply (-3/5) / (7/6) = -3/5 * 6/7 = - 18/35

OpenStudy (anonymous):

i totaly understand now thanks

OpenStudy (tal_tal):

18/35 sorry haven't done this in a while so if it's wrong I'm so sorry but how I do it is you flip the 7/6 to 6/7 then multiply

OpenStudy (texaschic101):

anytime :) But be sure and flip what you are dividing by and not the other number

OpenStudy (anonymous):

ok